OpenStudy (anonymous):

Which ordered pairs are solutions to the equation? y = x – 4 A. (10, 6) and (5, 1) B. (10, 6) and (1, 5) C. (10, 4) and (5, 1) D. (10, 5) and (5, 1)

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

@iYuko

OpenStudy (anonymous):

for each choices, you are given two different points. From these points, try plugging their x-values and solve for y. If you get the same value for y with the given point, then that will be the right choice. For example: the choice says (1,2) , x=1 then plug this into the equation: y= x - 4 y = 1-4 y=-3 y is not equal to 2, so this one isn't a solution to your equation. Try doing this with each point on your choices. Tell me which one satisfies the equation and I'll check it later ^_^

OpenStudy (anonymous):

i chose A

OpenStudy (anonymous):

:)